OpenStudy (anonymous):

2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (anonymous):

\[2x^2 - 2z^4 + y^2 - x^2 + z^4\]lets combine like terms first

OpenStudy (anonymous):

\(2x^2-x^2=x^2\) \(x^4-2z^4=-z^4\) so now it will be a bit easier if we write \[x^2+y^2-z^4\]

OpenStudy (anonymous):

make the replacements \[(-4)^2+3^2-(2)^4\] and arithmetic gets the answer

OpenStudy (haseeb96):

4-8+6+8+8 18 is answer